A gate of a reservoir is hinged at a point and it is prevented from opening by the force in the spring.
The gate is 3m high and 1m wide. The level of water is 2m above the top of the gate.
Determine the force in the spring if the gate is just about to open
1
Expert's answer
2021-08-30T01:55:08-0400

"Pressure = \rho gh\\
Pressure = 1000*9.8*2\\
\implies Force = Presure \times Area\\
Force = 1000*9.8*2 \times 3*1\\
\mathrm{Multiply\:the\:numbers:}\:1000\times \:9.8\times \:2\times \:3\times \:1=58800N"